OpenStudy (anonymous):

Solve each of the quadratic equations below and describe what the solution(s) represent to the graph of each. Show your work to receive full credit. y = x2 + 3x + 2 y = x2 + 2x + 1 Part 2: Using complete sentences, answer the following questions about the two quadratic equations above. Do the two quadratic equations have anything in common? If so, what? What makes y = x2 + 3x + 2 different from y = x2 + 2x + 1?

OpenStudy (btaylor):

so you want to set both equations equal to 0, then factor them. can you do that?

OpenStudy (anonymous):

@BTaylor thanks, yes i can do that

OpenStudy (btaylor):

ok, so you have it factored into: 1) (x+1)(x+2) = 0 2) (x+1)(x+1) = 0 right? So now you set both factors equal to 0: 1) x+1 = 0 AND x+2 = 0 so x = -1 and -2 2) x+1 = 0 AND x+1 = 0 so x = -1 They are different because in #2, it is a double root (you get x = -1 for both factors)

OpenStudy (btaylor):

Oh, and on the graph, the x-values are where they intersect the x-axis.

OpenStudy (anonymous):

click in my link

OpenStudy (anonymous):

I did @jcyogi

OpenStudy (anonymous):

and wud happen